OpenStudy (anonymous):

find number that arent the domain if the function f(y)=(2y-9)/(7y+6)

hartnn (hartnn):

the number for which the denominator of f(y) equals to 0 will NOT be in the domain. so, can you solve, 7y+6= 0 ?

OpenStudy (anonymous):

6/7?

hartnn (hartnn):

almost! try once again ? hint : be careful with signs.

OpenStudy (anonymous):

oh ok its negative 6/7

hartnn (hartnn):

thats correct, so x=-6/7 is the number not in the domain of f(y) thats it! any doubts ?

hartnn (hartnn):

** y= -6/7

OpenStudy (anonymous):

ok so just solve for the denominator to be undefined I got it

hartnn (hartnn):

good, well actually there are 2 things to consider in GENERAL, 1) denominator should NOT= 0 2) the quantity under the square root sign must be NON-NEGATIVE here there was no square root sign, so we just considered 1).

hartnn (hartnn):

and solve for denominator to be ZERO for which f(y) is undefined....